OpenStudy (vortish):

find the area of a triangle that has the height of 4/9 and a base of 1 3/4 and I know to find the area its one half of base times height i need help with the calculations

OpenStudy (accessdenied):

Well, I'd write it out like this first. \[ A = \frac{1}{2} b h \\ A = \frac{1}{2} \left( 1 \frac{3}{4} \right) \left( \frac{4}{9} \right) \] The mixed number isn't very useful here in multiplication. We should convert it into an improper fraction. \( 1\frac{3}{4} = \frac{3 + 4}{4} = \frac{7}{4} \) \[ A = \frac{1}{2} \left(\frac{7}{4} \right) \left( \frac{4}{9} \right) \\ A = \frac{1\times 7 \times 4}{2 \times 4 \times 9} \] Here, you may cancel the 4's in the numerator and denominator since it's essentially just 4/4= 1 Then just multiply out.

OpenStudy (vortish):

so you would come out with an equation it would be 1/2 x 7/9 right

OpenStudy (accessdenied):

A = 1/2 * 7/9 would be correct, although it would be considered proper to multiply it all together into one fraction.

OpenStudy (vortish):

7/18 or 14/9 i can not remember for the life of me right now sorry

OpenStudy (accessdenied):

For fractions, we just multiply straight across, numerator-to-numerator, denominator-to-denominator. :) So, 7/(2*9) = 7/18. :)

OpenStudy (vortish):

k thanks

OpenStudy (accessdenied):

You're welcome. :)